OpenStudy (anonymous):

MEDAL Class A of 20 students scored 60% while class B of 28 students undergoing the same test scored 70%. It was decided to put all marks together and find the common average

OpenStudy (anonymous):

(x1 + x2 + .. + x20)/20 = 0.60 (x1+x2 + ...+x20) = 19.6 (y1 + .. + y28)/28 = .70 (y1 + .. + y28) = 8.4 (x1 + .. + x20 + y1 + .. + y28) / (20+28) = (12+19.6)/48 = 65.83%

OpenStudy (anonymous):

oops I got a wrong number

OpenStudy (anonymous):

20*.6 = 12 and 28 * .7 = 19.6

OpenStudy (anonymous):

How did you get 19.6?

OpenStudy (anonymous):

Ok I get

OpenStudy (anonymous):

I corrected my numbers. The FINAL answer is right though

OpenStudy (anonymous):

Thanks for your help champ